a train leaves a station and travels east at 72 km/h. Three hours later a second train leaves on a parallel track and travels east at 120 km/h. When will it overtake the first train?
Answer by arunpaul(104)

let train A travels at speed = 72 kph
let train travels at speed = 120 kph
after 3 hours 2nd train i.e B train starts
so in 3 hour train a will cross = 3*72=216 km
the difference of speedof train A and train B = 120-72=48 kph
so to cover 216 km by the train B the time required = 216/48=4.5 hour
so after 4 hour and 30 minutes the train B will cross train A